# Microscopic Structure and Dynamics of Interfacial Water at Fluorinated vs Nonfluorinated Surfaces - Insights from Ab-Initio Simulations and IR Spectroscopy

### 27 Mar 2026 v1

@@ Abstract
Per- and polyfluoroalkyl substances are a class of synthetic chemical compounds widely used as coatings to lower surface energies. Yet the microscopic mechanisms of their weak interaction with water and organic compounds remain poorly understood. Here, we perform large-scale density-functional-theory molecular dynamics simulations to investigate water at self-assembled monolayers (SAMs) of fluorinated and non-fluorinated hydrocarbons. We analyze the interfacial water structure and compare it to the prototypical hydrophobic air–water interface. The interfacial water structure at both SAMs closely resembles that at the air–water interface, featuring a distinct depletion layer and a two-dimensional hydrogen-bond network parallel to the surface. Computed anisotropic infrared spectra reproduce key experimental signatures observed in surface-enhanced infrared absorption spectroscopy (SEIRAS), including the presence of free OH vibrations directly probing the local surface–water interactions. Notably, while the free OH stretch at the hydrocarbon SAM–water interface exhibits a red shift relative to the air–water interface, indicative of weak binding, the fluorinated SAM–water interface displays a weakly blue-shifted free OH mode, in agreement with experiment. This frequency behavior, which defies common interpretations based on the vibrational Stark effect, indicates that dispersive rather than electrostatic interactions dominate the interaction between water and SAMs. Analysis of spectral line shapes further shows that the reorientation dynamics of water molecules are significantly slower near the fluorinated surface, as commonly observed at hydrophilic surfaces. This indicates that fluorinated surfaces, despite being macroscopically more hydrophobic than their unfluorinated counterparts, exhibit spectroscopic characteristics that neither qualify it as hydrophobic nor hydrophilic.
@@ I Introduction
Per- and polyfluoroalkyl substances (PFAS) are a class of synthetic organic chemical compounds characterized by alkyl chains with multiple fluorine atoms attached. Due to their ability to modify the physical and chemical properties of various surfaces, these compounds hold significant scientific and technological importance in fields as diverse as nanoelectronics Klauk (2010); Xiang et al. (2016), biomolecular science Cametti et al. (2012); Lv and Cheng (2021), and electrochemistry Ozden et al. (2020). One of the most intriguing characteristics of PFAS is that the replacement of hydrogen atoms with fluorine atoms typically leads to an increase in hydrophobicity, as well as enhanced lipophobicity, a phenomenon known as omniphobicity. In the context of surface science, this means that PFAS-based surface coatings exhibit very low surface energies. Consequently, the incorporation of fluorochemicals is one of the primary techniques for achieving lower wettability of surfaces Tuteja et al. (2007); Si and Guo (2015); Parvate, Dixit, and Chattopadhyay (2020). Other characteristics that make PFAS-based coatings attractive include their high thermal and chemical stability Tuteja et al. (2007); Cametti et al. (2012); Soto et al. (2018). However, it has been found that the widespread use and long natural lifetimes of PFAS have led to their accumulation in the environment, and PFAS contamination has been linked to a variety of environmental problems and human diseases Brunn et al. (2023); Jian et al. (2018); Fenton et al. (2021). Thus, there is an ongoing effort to develop alternatives to PFAS-based coatings with comparably low surface energies. Understanding the microscopic mechanisms that govern the physicochemical characteristics of PFAS-based surface coatings can guide this effort.
@@
Experimentally, the microscopic structure of water at interfaces has been probed using various spectroscopic techniques, including IR and THz absorption spectroscopy Le Caër et al. (2011); Knight et al. (2019); Pezzotti et al. (2021); Ruiz-Barragan et al. (2022); Zhang et al. (2025), and surface-specific nonlinear spectroscopic methods, such as sum frequency generation (SFG) Du et al. (1993); Du, Freysz, and Shen (1994); Perakis et al. (2016); Gonella et al. (2021) . Different spectroscopic fingerprints of water at hydrophobic interfaces have been established. Particular well studied is the emergence of so-called ’dangling’ or ’free’ OH bonds Du et al. (1993); Du, Freysz, and Shen (1994); Perakis et al. (2016); Pezzotti et al. (2021), corresponding to water molecules in the outermost layer, whose OH bonds point towards the hydrophobic interface and are therefore not hydrogen bonded. These water molecules give rise to distinct spectroscopic markers in SFG spectra: the emergence of a narrow band at around 3700 cm-1, a frequency higher than the typical bulk OH stretch peak. Signatures of the same population of water molecules have also been analyzed using x-ray absorption spectroscopy Velasco-Velez et al. (2014) and surface enhanced infrared absorption spectroscopy Ataka, Yotsuyanagi, and Osawa (1996). Since spectroscopic signatures of free OH groups are typically found at hydrophobic interfaces, the intensity of the corresponding band has been proposed as a microscopic marker of hydrophobicity Tang et al. (2020). However, recent experiments have shown that free OH groups can also be found at macroscopically hydrophilic surfaces Tuladhar, Piontek, and Borguet (2017); Cyran et al. (2019); Wang et al. (2019). Instead, the vibrational frequency of free OH groups measured in sum frequency generation spectroscopy has been proposed as an indicator of hydrophobicity Pezzotti et al. (2021). Comparing SFG spectra of water at a range of different surfaces, it was shown that the spectral peak corresponding to the vibration of free OH groups shifts to lower frequencies at more hydrophilic interfaces. This can be interpreted within the framework of the vibrational Stark effect: Stronger electrostatic interactions between water molecules and surfaces, which give rise to the increased hydrophilicity of the surface, lead to a red-shift of the free OH vibrational feature Auer and Skinner (2008); Bishop (1993). A different framework for interpreting vibrational frequency shifts was recently introduced by Brünig et al, in which vibrational blue shifts were related to frequency-dependent friction Brünig et al. (2022).
@@
In this work, we report on large-scale density-functional-theory-based (DFT) MD simulations of the SAM-water interface for perfluoroalkylthiol-SAMs and alkanethiol-SAMs. We additionally compare our simulation results to the prototypical hydrophobic interface, the air-water interface. Our findings demonstrate that the microscopic water structure predicted by DFT-MD aligns well with results from empirical force field simulations. Moreover, key microscopic descriptors of the surface’s hydrophobicity, such as the thickness of the interfacial depletion layer, show qualitative consistency with both experimental data and previous computational predictions. The water’s hydrogen bond network in the vicinity of all three interfaces is remarkably similar: we observe the formation of a layer of water molecules that are predominantly hydrogen bonded with other water molecules in the same layer, as was previously reported for various hydrophobic surfaces Pezzotti, Galimberti, and Gaigeot (2017).
@@ II Results and discussion
We conduct DFT-MD simulations of the HSAM-water interface (Fig. 1 B), comprising 16 n-octane molecules (Fig. 1 D) and 198 water molecules as well as of the FSAM-water interface (Fig. 1 C) comprising 16 semifluorinated molecules CH3CH2(CF2)3CF3 (Fig. 1 D) and 286 water molecules. Carbon atoms at the bottom of the SAMs are restrained to a hexagonal grid (Fig. 1 D-E) using harmonic restraint potentials. The lattice constants are chosen to align with the experimentally determined values of a = 4.97 for HSAMs and 5.9 for FSAMs Chidsey and Loiacono (1990); Alves and Porter (1993); Barriet and Lee (2003). We compare our DFT-MD simulations to FF-MD simulations comprising 200 nonfluorinated and semifluorinated hexane molecules each, and 2089 and 2928 water molecules, respectively. Water molecules are modeled using the SPC/E model Berendsen, Grigera, and Straatsma (1987), while hydrocarbons are modeled using the OPLS/AA FF Jorgensen, Maxwell, and Tirado-Rives (1996) and its extension for perfluorinated hydrocarbons Watkins and Jorgensen (2001). In a previous work, we have adapted this FF to contain parameters for semifluorinated hydrocarbons Carlson et al. (2021). In the same work, we have shown that the employed FFs can quantitatively reproduce the experimental contact angles of water droplets on both HSAMs and FSAMs with high accuracy. The primary mechanism responsible for the increased hydrophobicity of FSAMs, relative to HSAMs, was identified as the larger lattice spacing of semifluorinated monolayers Carlson et al. (2021). Further simulation details can be found in the Methods section and snapshots of the full simulation setups are shown in Sec. S1 of the supplementary material.
The prototypical hydrophobic interface is the air-water interface. To identify surface-specificity of our results, we compare to a previously publishedBecker, Loche, and Netz (2022) DFT-MD simulation of 352 water molecules within a periodic box of size nm3. These simulations were conducted using the same exchange-correlation functional and simulation parameters, rendering them comparable to the SAM-water interface simulations reported in this work. The same system was also simulated using FF-MD simulations employing the SPC/E force field Berendsen, Grigera, and Straatsma (1987). All details of the air-water simulations are provided in Becker, Loche, and Netz (2022).
@@ II.1 Structure of interfacial water
Experimentally, a surface’s hydrophobicity is typically characterized by the contact angle that a water droplet forms on it. While the accurate determination of contact angles via FF-MD simulations is feasible, it requires either the simulation of very large systems when explicit water droplets are considered Carlson et al. (2021); De Ruijter, Blake, and De Coninck (1999); Kanduč (2017) or very long simulation times when surface energies are estimated from the anisotropy of the pressure tensor Carlson et al. (2024). Both methods are prohibitively expensive for implementation with DFT-MD simulations. Instead, to establish the hydrophobicity of HSAMs and FSAMs through DFT-MD simulations, we rely on comparing microscopic observables that characterize surface energy between our DFT-MD simulations and FF-MD simulations for which contact angles have previously been established Carlson et al. (2021). For instance it has been shown that the density depletion between a solid surface and water is strongly correlated to the contact angle Janecek and Netz (2007); Huang et al. (2008); Sendner et al. (2009); Carlson and Netz (2025). Since mass density profiles converge already for short simulation times, this depletion is a good observable to compare the surface energy of different surfaces from DFT-MD simulations.
@@
to determine the Gibbs’s dividing surface position of the liquid, can be calculated according to Mamatkulov, Khabibullaev, and Netz (2004)
@@
After establishing that our DFT-MD simulations reproduce the interfacial properties found in experimental and previous theoretical work, we analyze how this relates to the microscopic water structure in the vicinity of the interface, particularly in comparison with the prototypical hydrophobic interface: the air-water interface. It is important to note that, due to interfacial nano roughness and the occurrence of capillary waves at aqueous interfaces, the profiles of interfacial properties depend on the scale at which they are observed Sedlmeier, Horinek, and Netz (2009). To mitigate the influence of varying surface roughness on the comparison between the air-water and SAM-water interfaces, we calculate the Willard-Chandler interface (WCI), which constructs an instantaneous and locally flexible interface with respect to which interfacial properties can be compared Willard and Chandler (2010). Indeed, when determined in the laboratory frame, the water density profile at the air-water interface is fairly broad and approximately sigmoidal in shape (Fig. 1 G), while a stratified density is observed for the SAM-water interface (Fig. 1 H-I). In contrast, when determined with respect to the WCI, the molecular number density profiles of all three interfaces reveal remarkably similar stratification (Fig. 2 A-C), indicating that the microscopic molecular structure of water at alkane SAMs is primarily governed by water-water interactions rather than water-surface interactions. For instance, the position of the first minimum of the density profile, which in our notation defines the first hydration layer (FHL) of the interface, is found to be identical for all three interfaces at . Additionally, the density maxima are of similar magnitude. In Sec. S3 of the supplementary material we shows that the same density stratification is found from FF-MD simulations, corroborating our conclusion that DFT-MD and FF-MD simulations predict similar water-surface interactions.
@@
To study the arrangement of water in the first hydration layer of the interface, we perform orientational analysis of water molecules. The orientation of a water molecule can be characterized by the two angles and , which correspond to the angles that the molecule’s dipole vector and the angle an OH bond form with the interfacial normal, respectively Sedlmeier et al. (2008). The joint orientational distribution functions in Fig. 2 D,F,H, reveal a strongly anisotropic orientation of water molecules in the outermost layer of water () at all three interfaces. In this region, consistent with the observations from the density profiles depicted in Figs. 2 A-C, the orientation distribution functions further demonstrate that the structure of interfacial water is remarkably similar across all three interfaces considered. Regarding orientation, we find two major populations of water molecules: The first population consists of molecules in which one of the OH bonds points away from the bulk phase toward the interface (), while the other OH bond is oriented toward the bulk phase. This population corresponds to the two probability maxima in Figs. 2D-F around () and (). Corresponding molecular orientations are indicated in Fig. 2 J. Note that, due to the fact that each water molecule has two OH bonds, both maxima in the probability distribution correspond to the same population of water molecules. The second population that can be found in the outermost layer of the interface, consists of water molecules with an in-plane orientation, whose OH bonds point only slightly toward the bulk phase (the corresponding maximum is located around () and the corresponding orientation is shown in Fig. 2 J). The orientation of water molecules in the remainder of the first hydration layer in the region is shown in Fig. 2 G-I. Overall, we observe less orientational anisotropy compared to the immediate vicinity of the interface. The main feature is an increased tendency for molecules to align their OH bonds parallel to the interface. These results are consistent with previous experimental and theoretical work Sedlmeier et al. (2008); Pezzotti, Galimberti, and Gaigeot (2017); Pezzotti, Serva, and Gaigeot (2018); Pezzotti et al. (2021); Lehmann et al. (2025) that found that at hydrophobic interfaces, water forms a layer of molecules, which are strongly connected to each other via in-plane hydrogen bonds.

### 03 Apr 2026 v2

@@ Abstract
Per- and polyfluoroalkyl substances are a class of synthetic chemical compounds widely used as coatings to lower surface energies. Yet the microscopic mechanisms of their weak interaction with water and organic compounds remain poorly understood. Here, we perform large-scale density-functional-theory molecular dynamics simulations to investigate water at self-assembled monolayers (SAMs) of fluorinated and non-fluorinated hydrocarbons. We analyze the interfacial water structure and compare it to the prototypical hydrophobic air–water interface. The interfacial water structure at both SAMs closely resembles that at the air–water interface, featuring a distinct depletion layer and a two-dimensional hydrogen-bond network parallel to the surface. Computed anisotropic infrared spectra reproduce key experimental signatures observed in surface-enhanced infrared absorption spectroscopy (SEIRAS), including the presence of free OH vibrations directly probing the local surface–water interactions. Notably, while the free OH stretch at the hydrocarbon SAM–water interface exhibits a red shift relative to the air–water interface, indicative of weak binding, the fluorinated SAM–water interface displays a weakly blue-shifted free OH mode, in agreement with experiment. This frequency behavior defies common interpretations based on the vibrational Stark effect. Further, we show that the reorientation dynamics of water molecules are significantly slower near the fluorinated surface, which we correlate with spectral line shapes, an effect rather expected at hydrophilic surfaces. This indicates that fluorinated SAMs, despite being macroscopically more hydrophobic than their unfluorinated counterparts, exhibit spectroscopic characteristics that neither qualify it as hydrophobic nor hydrophilic.
@@ I Introduction
Per- and polyfluoroalkyl substances (PFAS) are a class of synthetic organic chemical compounds characterized by alkyl chains with multiple fluorine atoms attached. Due to their ability to modify the physical and chemical properties of various surfaces, these compounds hold significant scientific and technological importance in fields as diverse as nanoelectronics Klauk (2010); Xiang et al. (2016), biomolecular science Cametti et al. (2012); Lv and Cheng (2021), and electrochemistry Ozden et al. (2020). One of the most intriguing characteristics of PFAS is that the replacement of hydrogen atoms with fluorine atoms typically leads to an increase in hydrophobicity, as well as enhanced lipophobicity, a phenomenon known as omniphobicity. In the context of surface science, this means that PFAS-based surface coatings exhibit very low surface energies. Consequently, the incorporation of fluorochemicals is one of the primary techniques for achieving lower wettability of surfaces Tuteja et al. (2007); Si and Guo (2015); Parvate et al. (2020). Other characteristics that make PFAS-based coatings attractive include their high thermal and chemical stability Tuteja et al. (2007); Cametti et al. (2012); Soto et al. (2018). However, it has been found that the widespread use and long natural lifetimes of PFAS have led to their accumulation in the environment, and PFAS contamination has been linked to a variety of environmental problems and human diseases Brunn et al. (2023); Jian et al. (2018); Fenton et al. (2021). Thus, there is an ongoing effort to develop alternatives to PFAS-based coatings with comparably low surface energies. Understanding the microscopic mechanisms that govern the physicochemical characteristics of PFAS-based surface coatings can guide this effort.
@@
Experimentally, the microscopic structure of water at interfaces has been probed using various spectroscopic techniques, including IR and THz absorption spectroscopy Le Caër et al. (2011); Knight et al. (2019); Pezzotti et al. (2021); Ruiz-Barragan et al. (2022); Zhang et al. (2025), and surface-specific nonlinear spectroscopic methods, such as sum frequency generation (SFG) Du et al. (1993, 1994); Perakis et al. (2016); Gonella et al. (2021) . Different spectroscopic fingerprints of water at hydrophobic interfaces have been established. Particular well studied is the emergence of so-called ’dangling’ or ’free’ OH bonds Du et al. (1993, 1994); Perakis et al. (2016); Pezzotti et al. (2021), corresponding to water molecules in the outermost layer, whose OH bonds point towards the hydrophobic interface and are therefore not hydrogen bonded. These water molecules give rise to distinct spectroscopic markers in SFG spectra: the emergence of a narrow band at around 3700 cm-1, a frequency higher than the typical bulk OH stretch peak. Signatures of the same population of water molecules have also been analyzed using x-ray absorption spectroscopy Velasco-Velez et al. (2014) and surface enhanced infrared absorption spectroscopy Ataka et al. (1996). Since spectroscopic signatures of free OH groups are typically found at hydrophobic interfaces, the intensity of the corresponding band has been proposed as a microscopic marker of hydrophobicity Tang et al. (2020). However, recent experiments have shown that free OH groups can also be found at macroscopically hydrophilic surfaces Tuladhar et al. (2017); Cyran et al. (2019); Wang et al. (2019). Instead, the vibrational frequency of free OH groups measured in sum frequency generation spectroscopy has been proposed as an indicator of hydrophobicity Pezzotti et al. (2021). Comparing SFG spectra of water at a range of different surfaces, it was shown that the spectral peak corresponding to the vibration of free OH groups shifts to lower frequencies at more hydrophilic interfaces. This can be interpreted within the framework of the vibrational Stark effect: Stronger electrostatic interactions between water molecules and surfaces, which give rise to the increased hydrophilicity of the surface, lead to a red-shift of the free OH vibrational feature Auer and Skinner (2008); Bishop (1993). A different framework for interpreting vibrational frequency shifts was recently introduced by Brünig et al, in which vibrational blue shifts were related to frequency-dependent friction Brünig et al. (2022).
@@
In this work, we report on large-scale density-functional-theory-based (DFT) MD simulations of the SAM-water interface for perfluoroalkylthiol-SAMs and alkanethiol-SAMs. We additionally compare our simulation results to the prototypical hydrophobic interface, the air-water interface. Our findings demonstrate that the microscopic water structure predicted by DFT-MD aligns well with results from empirical force field simulations. Moreover, key microscopic descriptors of the surface’s hydrophobicity, such as the thickness of the interfacial depletion layer, show qualitative consistency with both experimental data and previous computational predictions. The water’s hydrogen bond network in the vicinity of all three interfaces is remarkably similar: we observe the formation of a layer of water molecules that are predominantly hydrogen bonded with other water molecules in the same layer, as was previously reported for various hydrophobic surfaces Pezzotti et al. (2017).
@@ II Results and discussion
We conduct DFT-MD simulations of the HSAM-water interface (Fig. 1 B), comprising 16 n-octane molecules (Fig. 1 D) and 198 water molecules as well as of the FSAM-water interface (Fig. 1 C) comprising 16 semifluorinated molecules CH3CH2(CF2)3CF3 (Fig. 1 D) and 286 water molecules. Carbon atoms at the bottom of the SAMs are restrained to a hexagonal grid (Fig. 1 D-E) using harmonic restraint potentials. The lattice constants are chosen to align with the experimentally determined values of a = 4.97 for HSAMs and 5.9 for FSAMs Chidsey and Loiacono (1990); Alves and Porter (1993); Barriet and Lee (2003). We compare our DFT-MD simulations to FF-MD simulations comprising 200 nonfluorinated and semifluorinated hexane molecules each, and 2089 and 2928 water molecules, respectively. Water molecules are modeled using the SPC/E model Berendsen et al. (1987), while hydrocarbons are modeled using the OPLS/AA FF Jorgensen et al. (1996) and its extension for perfluorinated hydrocarbons Watkins and Jorgensen (2001). In a previous work, we have adapted this FF to contain parameters for semifluorinated hydrocarbons Carlson et al. (2021). In the same work, we have shown that the employed FFs can quantitatively reproduce the experimental contact angles of water droplets on both HSAMs and FSAMs with high accuracy. The primary mechanism responsible for the increased hydrophobicity of FSAMs, relative to HSAMs, was identified as the larger lattice spacing of semifluorinated monolayers Carlson et al. (2021). Further simulation details can be found in the Methods section and snapshots of the full simulation setups are shown in Sec. S1 of the supplementary material.
The prototypical hydrophobic interface is the air-water interface. To identify surface-specificity of our results, we compare to a previously publishedBecker et al. (2022) DFT-MD simulation of 352 water molecules within a periodic box of size nm3. These simulations were conducted using the same exchange-correlation functional and simulation parameters, rendering them comparable to the SAM-water interface simulations reported in this work. The same system was also simulated using FF-MD simulations employing the SPC/E force field Berendsen et al. (1987). All details of the air-water simulations are provided in Becker et al. (2022).
@@ II.1 Structure of interfacial water
Experimentally, a surface’s hydrophobicity is typically characterized by the contact angle that a water droplet forms on it. While the accurate determination of contact angles via FF-MD simulations is feasible, it requires either the simulation of very large systems when explicit water droplets are considered Carlson et al. (2021); De Ruijter et al. (1999); Kanduč (2017) or very long simulation times when surface energies are estimated from the anisotropy of the pressure tensor Carlson et al. (2024). Both methods are prohibitively expensive for implementation with DFT-MD simulations. Instead, to establish the hydrophobicity of HSAMs and FSAMs through DFT-MD simulations, we rely on comparing microscopic observables that characterize surface energy between our DFT-MD simulations and FF-MD simulations for which contact angles have previously been established Carlson et al. (2021). For instance it has been shown that the density depletion between a solid surface and water is strongly correlated to the contact angle Janecek and Netz (2007); Huang et al. (2008); Sendner et al. (2009); Carlson and Netz (2025). Since mass density profiles converge already for short simulation times, this depletion is a good observable to compare the surface energy of different surfaces from DFT-MD simulations.
@@
to determine the Gibbs’s dividing surface position of the liquid, can be calculated according to Mamatkulov et al. (2004)
@@
After establishing that our DFT-MD simulations reproduce the interfacial properties found in experimental and previous theoretical work, we analyze how this relates to the microscopic water structure in the vicinity of the interface, particularly in comparison with the prototypical hydrophobic interface: the air-water interface. It is important to note that, due to interfacial nano roughness and the occurrence of capillary waves at aqueous interfaces, the profiles of interfacial properties depend on the scale at which they are observed Sedlmeier et al. (2009). To mitigate the influence of varying surface roughness on the comparison between the air-water and SAM-water interfaces, we calculate the Willard-Chandler interface (WCI), which constructs an instantaneous and locally flexible interface with respect to which interfacial properties can be compared Willard and Chandler (2010). Indeed, when determined in the laboratory frame, the water density profile at the air-water interface is fairly broad and approximately sigmoidal in shape (Fig. 1 G), while a stratified density is observed for the SAM-water interface (Fig. 1 H-I). In contrast, when determined with respect to the WCI, the molecular number density profiles of all three interfaces reveal remarkably similar stratification (Fig. 2 A-C), indicating that the microscopic molecular structure of water at alkane SAMs is primarily governed by water-water interactions rather than water-surface interactions. For instance, the position of the first minimum of the density profile, which in our notation defines the first hydration layer (FHL) of the interface, is found to be identical for all three interfaces at . Additionally, the density maxima are of similar magnitude. In Sec. S3 of the supplementary material we shows that the same density stratification is found from FF-MD simulations, corroborating our conclusion that DFT-MD and FF-MD simulations predict similar water-surface interactions.
@@
To study the arrangement of water in the first hydration layer of the interface, we perform orientational analysis of water molecules. The orientation of a water molecule can be characterized by the two angles and , which correspond to the angles that the molecule’s dipole vector and the angle an OH bond form with the interfacial normal, respectively Sedlmeier et al. (2008). The joint orientational distribution functions in Fig. 2 D,F,H, reveal a strongly anisotropic orientation of water molecules in the outermost layer of water () at all three interfaces. In this region, consistent with the observations from the density profiles depicted in Figs. 2 A-C, the orientation distribution functions further demonstrate that the structure of interfacial water is remarkably similar across all three interfaces considered. Regarding orientation, we find two major populations of water molecules: The first population consists of molecules in which one of the OH bonds points away from the bulk phase toward the interface (), while the other OH bond is oriented toward the bulk phase. This population corresponds to the two probability maxima in Figs. 2D-F around () and (). Corresponding molecular orientations are indicated in Fig. 2 J. Note that, due to the fact that each water molecule has two OH bonds, both maxima in the probability distribution correspond to the same population of water molecules. The second population that can be found in the outermost layer of the interface, consists of water molecules with an in-plane orientation, whose OH bonds point only slightly toward the bulk phase (the corresponding maximum is located around () and the corresponding orientation is shown in Fig. 2 J). The orientation of water molecules in the remainder of the first hydration layer in the region is shown in Fig. 2 G-I. Overall, we observe less orientational anisotropy compared to the immediate vicinity of the interface. The main feature is an increased tendency for molecules to align their OH bonds parallel to the interface. These results are consistent with previous experimental and theoretical work Sedlmeier et al. (2008); Pezzotti et al. (2017, 2018, 2021); Lehmann et al. (2025) that found that at hydrophobic interfaces, water forms a layer of molecules, which are strongly connected to each other via in-plane hydrogen bonds.